OpenStudy (anonymous):

If 18 identical machines required 40 days to complete a job, how many fewer days would have been required to do the job if 6 additional machines of the same type had been used from the beginning?

OpenStudy (anonymous):

each machine has a rate of 1/(18*40). 6 more machines added, then 1/(18*40) = 1/(24 d) solve for d and d = 30 days. Which means it will take 10 fewer days

OpenStudy (anonymous):

am not getting you.. any simple step...

OpenStudy (anonymous):

well, what if 18 machines can finish the job in 40 days, then, in 40 days, each machine can do 1/18 job yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it now... :-) thanks...

OpenStudy (anonymous):

ok, i'll finish: then in 1 day, each machine can do 1/(18 *40) job then in 1 day, 26 machines can do 26/(18*40) job which means it will take 26 machines 1/[26/(18*40)] = 30 days day, which means 10 days fewer

OpenStudy (anonymous):

bundle of thanks dear... :-)